OpenStudy (anonymous):

Look at the circle shown below. What is the equation of the circle? Answer x2 + (y – 2)2 = 36 x2 + (y – 2)2 = 6 (x – 2)2 + y 2 = 36 (x – 2)2 + y 2 = 6

Parth (parthkohli):

@lalaly =)

OpenStudy (anonymous):

no circle to look.....

Parth (parthkohli):

Anyway: \( \color{Black}{\Rightarrow (x - h)^2 + ( y - k)^2 = r^2 }\) See the radius and then the center coordinates .

Parth (parthkohli):

Center is (h,k)

Parth (parthkohli):

I see that the center is \((0,2)\) here.

OpenStudy (anonymous):

r is radius

Parth (parthkohli):

\( \color{Black}{\Rightarrow (x - 0)^2 + (y - 2)^2 = 6^2 }\)

Parth (parthkohli):

^ you can simplify that

OpenStudy (anonymous):

it looks like the first opton but do i keep the 6 or do i work it out and change t to 36

OpenStudy (anonymous):

still need help?

Parth (parthkohli):

What is 6^2?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

\[(x-2)^{2}+y ^{2}=36\]